# PCL-Indexability and Whittle Index for Restless Bandits with General Observation Models

### v3

@@ 3.3 Numerical Results
In this subsection, we present the robust performance of our algorithm through extensive numerical simulations. To approximate the state space, we set the maximum number of iterations to 6 and the error limit $\varepsilon$ to $31e-$ 3. We expanded our simulation framework to evaluate systems of varying scales, ranging from 3x3 to 6x6 configurations (for the transition matrix of arm states). For each arm state size, we conducted simulations with 30, 40, and 50 arms, while maintaining the activation of a single arm per time slot and system parameters adhered to the PCL-indexability. Our primary focus is on comparing the unit-time-return performance of the PCL/Whittle index policy against a myopic policy over 200 time slots, utilizing 10,000 Monte Carlo simulation experiments for each configuration. The performance gap, quantified as the percentage of improvement from the Whittle index policy over the myopic policy, is summarized in Table 1. As shown in all the figures below, the PCL/Whittle index policy consistently achieves significantly superior performance compared to the myopic policy.
@@
4 Conclusion
@@
In this paper, we generalized the PCL framework for RMAB with finite state spaces to countable state spaces and propose an efficient algorithm to test the PCL-indexability and compute the Whittle index function for the partially observable RMAB with general observation models. As the number of belief states increases rapidly with time, the classic dynamic programming analysis for the RMAB problem suffers from the curse of dimensionality. Based on the achievable region and conservation law analysis studied in the literature for various RMAB problems, this paper is the first to extend the previous theoretical results in this direction to infinite-state scenarios by taking advantages of the theory of infinite-dimensional linear programming. Under a loose regularity assumption, we prove that the PCL-indexability is a sufficient condition for the Whittle indexability. Last, by introducing the concept of approximate state space, we build an algorithm to calculate the approximate Whittle index of each state using a finite-state adaptive greedy algorithm. Numerical results verified the strong performance of the Whittle index policy in our model.
@@
Although we have quite a few theoretical analyses and have confirmed the effectiveness of the Whittle index algorithm, some problems remain to be solved in the future. An important one is that we only analyzed the optimality of the Whittle index under the condition that the problem satisfies PCL-indexability with respect to a specific set family $\tilde{\Omega}$, but we have not yet explored other possible sets that might also allow PCL to function effectively. Furthermore, the strict calculation of Whittle indices for the entire belief state space remains an open problem (even though only their relative order is required for optimal policy implementation). In summary, there are still many opportunities for further exploration of problems with infinite state spaces.

### v4

@@ 3.3 Numerical Results
In this subsection, we present the robust performance of our algorithm through extensive numerical simulations. To approximate the state space, we set the maximum number of iterations to 6 and the error limit $\varepsilon$ to $10^{-3}$. We expanded our simulation framework to evaluate systems of varying scales, ranging from 3x3 to 6x6 configurations (for the transition matrix of arm states). For each arm state size, we conducted simulations with 30, 40, and 50 arms, while maintaining the arm selection and system parameters adhered to the PCL-indexability. Our primary focus is on comparing the unit-time-return performance of the PCL/Whittle index policy against a myopic policy over 200 decision epochs, utilizing 10,000 Monte Carlo simulation experiments for each configuration. The performance gap, quantified as the percentage of improvement from the Whittle index policy over the myopic policy, is summarized in Table 2. As shown in Fig. 2-5, the PCL/Whittle index policy consistently achieves significantly superior performance compared to the myopic policy.
@@
Next, we compare our Whittle index policy with more baselines to demonstrate the strong performance of the former.
(a.1) Rollout and VFA baselines: We now compare our Whittle index policy with (i) a one-step rollout policy based on the myopic baseline [1], and (ii) value-function approximation policies (QMDP and FIB) [2], across four discount factors $\beta\in\{0.7,0.8,0.9,0.9999\}$. Figure 6 plots the average reward over time for all five policies.
Figure 6: Performance of Whittle index, myopic, rollout, QMDP-VFA, and FIB-VFA policies.
Across all four discount factors, the Whittle index policy consistently outperforms the rollout baseline. The two VFA policies closely track the myopic policy in both performance and runtime (see Figure 7); their curves are almost indistinguishable from myopic in Figure 6. Rollout is far more computationally expensive than Whittle, without providing a clear performance advantage,and the two VFA baselines perform nearly identically to the myopic policy, we therefore omit them from the main (large-scale) experiments to maintain figure clarity. The myopic policy alone serves as a consolidated proxy for these simple baselines.
Runtime versus number of arms and time horizon: On the experiment above, we measure the computation time per simulation run as a function of number of arms $N$ or time horizon $T$, for all five policies (myopic, Whittle, rollout, QMDP-VFA, and FIB-VFA). Figure 7 shows the resulting curves with a logarithmic $y$-axis.
Figure 7: Computation time versus NN and $T$.
The myopic, Whittle index, and VFA policies all exhibit roughly linear growth in $N$ and $T$, whereas the rollout baseline becomes several orders of magnitude slower as $N$ or $T$ increases. This confirms that our Whittle index computation scales well in terms of $N$ and $T$, while rollout is not practical beyond small instances.
(a.2) Exact dynamic programming on tiny instances: To compare against the exact dynamic-programming, we consider a tiny two-state, four-arm instance with time horizon $7$, for which dynamic programming is feasible. We generate four independent instances, for each one, compute both the optimal (dynamic-programming) policy and the Whittle policy. Figure 8 plots the average discounted reward over time for the two policies, based on $500$ Monte–Carlo runs per instance.
Figure 8: Performance of the optimal and Whittle index policies.
On three of the four instances (panels 1, 3, and 4), the Whittle and optimal curves coincide at all time steps up to numerical precision. On the remaining instance (panel 2), the Whittle policy starts slightly below the optimal policy in the very early steps (with a maximum percentage gap of about $0.24\%$ in average reward), but quickly catches up; by the final time step the Whittle policy lags the optimal one by only $0.01\%$.
Thus, whenever dynamic programming is computationally feasible, the Whittle policy essentially recovers the optimal performance trajectory.
However, computing the exact dynamic-programming policy is already very expensive on this tiny example: with 500 Monte–Carlo runs, the average runtime is $61.58$ s and the peak resident set size (RSS) is $1768.40$ MB. By contrast, our Whittle policy on the larger three-state, six-arm benchmark runs in a fraction of a second per arm and uses under $100$ MB of memory (see part (b) below). This confirms that dynamic programming is useful as a sanity check on small instances, but it cannot serve as a scalable baseline for large systems.
(b) Runtime and memory
Memory usage: We monitor peak RSS on a three-state, six-arm instance with time horizon $100$ and $500$ Monte–Carlo runs. Table 3 reports the measured peak RSS for each policy. All five policies have essentially the same peak memory footprint.
Table 3: Peak resident set size.
Policy Peak RSS (MB)
Myopic 92.75
Whittle 93.14
Rollout (1-step) 93.72
QMDP-VFA 94.05
FIB-VFA 94.10
[…]
Figure 9: Performance of Whittle index and myopic policies on larger $K\times K$ instances.
Table 4: Performance gap and runtime on larger $K\times K$ instances.
State dimension $K$ Whittle improvement over myopic (%) runtime (s)
6 8.90 1820.96
7 7.75 2139.01
8 6.01 2862.12
9 6.12 2993.83
Now we study the performance of our Whittle index policy under four observation-error structures (symmetric, state-independent, asymmetric, and random matrices). For each structure, we record the average index computation time per arm, and then run the Monte–Carlo simulation to compare the long-run performance of the Whittle index policy with the Myopic policy.
Figure 10: Whittle versus Myopic performance to four observation-error structures.
The index computation times are very similar across all four cases: the average time per arm ranges from $0.2278$ s to $0.2332$ s. In terms of policy performance, the Whittle index policy consistently dominates the myopic policy under all four observation models; Table 5 reports the improvement and computation times. These experiments show that (i) the index computation cost is essentially unchanged across a wide range of observation-error structures, and (ii) the Whittle index policy performs better than the myopic policy under any type of observation matrices.
Table 5: Performance under different observation matrices
Error structure Improvement (%) Computation time (s)
Symmetric 6.17 0.2332
State-independent 6.67 0.2317
Asymmetric 6.21 0.2290
Random 6.69 0.2278
Finally, we link the choice of $(T,\epsilon)$ to both the index error and the long-run performance, using a three-state, six-arm instance with fixed transition, observation, and reward matrices and $1000$ times Monte Carlo simulations.
(a) Index error versus $(T,\epsilon)$. We first fix the parameters $(\bar{T},\bar{\epsilon})$ and use the approximate Whittle index $\bar{W}=W(\bar{T},\bar{\epsilon})$ as an almost exact index, with
$\bar{T}=50,\qquad\bar{\epsilon}=10^{-6}.$
We then vary $T$ and $\epsilon$ and record the absolute error $\bigl|W(T,\epsilon)-\bar{W}\bigr|$. The two plots are shown in Figure 11.
Figure 11: Left: index error versus $T$ (with $\epsilon=10^{-3}$). Right: index error versus $\epsilon$ (with $T=8$).
  • $T$: The index error decreases from $2.65\times 10^{-2}$ at $T=4$ to $1.02\times 10^{-2}$ at $T=10$, and then levels off at $3.28\times 10^{-3}$ for all $T\geq 17$.
  • $\epsilon$: As we decrease $\epsilon$ from $10^{-2}$ to $2\times 10^{-3}$, the error drops from $4.93\times 10^{-2}$ to $1.93\times 10^{-2}$; further tightening $\epsilon$ to $10^{-3}$ and $5\times 10^{-4}$ only reduce the error slightly to about $1.40\times 10^{-2}$.
These results show that the $T$ is the main factor controlling the index error: once $T$ is moderately large, the additional gain is small. Reducing $\epsilon$ from a coarse value ($\approx 10^{-2}$) to a moderately small value (at the order of $\times 10^{-3}$) removes most of the error, and further reductions have limited effect. Since the error curve in $\epsilon$ already flattens out around $\epsilon=10^{-3}$, we use this value in the more complex experiments and focus on the sensitivity to $T$.
(b) Impact on policy performance. To check the impact of $T$ on the policy performance, we keep the same three-state, six-arm instance and compare the long-run performance of the Whittle index policy (with different approximation depths $T$) to the myopic policy, under the same observation and reward model and with $\epsilon=10^{-3}$. The results are shown in Figure 12.
Figure 12: Performance of Whittle index and the myopic policies for different $T\in\{6,8,10,20\}$.
At the end of the simulation, the approximate Whittle index policy consistently dominates the myopic policy: the relative improvement in long-run average reward is about $9\%\text{--}10\%$, indicating that varying $T$ in this range has little effect on the final performance. Together with the index-error analysis above, the results indicate that setting $T$ around $6$ to $8$ achieves an effective balance between computational cost and policy performance for this problem instance.
4 Conclusion and Future Work
@@
In this paper, we generalized the PCL framework for RMAB with finite state spaces to countable state spaces and propose an efficient algorithm to test the PCL-indexability and compute the Whittle index function for the partially observable RMAB with general observation models. As the number of belief states increases rapidly with time, the classic dynamic programming analysis for the RMAB problem suffers from the curse of dimensionality. Based on the achievable region and conservation law analysis studied in the literature for various RMAB problems, this paper is the first to extend the previous theoretical results in this direction to high-dimensional infinite-state scenarios by taking advantages of the theory of infinite-dimensional linear programming. Under a loose regularity assumption, we prove that the PCL-indexability is a sufficient condition for the Whittle indexability. Last, by introducing the concept of approximate state space, we build an algorithm to calculate the approximate Whittle index of each state using a finite-state adaptive greedy algorithm. Numerical results verified the strong performance of the Whittle index policy in our model.
@@
Although we have quite a few theoretical analyses and have confirmed the effectiveness of the Whittle index algorithm, some problems remain to be solved in the future. An important one is that we only analyzed the optimality of the Whittle index under the condition that the problem satisfies PCL-indexability with respect to a specific set family $\tilde{\Omega}$, but we have not yet explored other possible sets that might also allow PCL to function effectively. Furthermore, the strict calculation of Whittle indices for the entire belief state space remains an open problem (even though only their relative order is required for optimal policy implementation). Future work also includes the establishment of sufficient conditions for PCL-indexability and the extension of fluid analysis to the high-dimensional belief state space (bertsimas2016fluid; verloop2016rmab). In summary, there are still many opportunities for further exploration of problems with infinite state spaces.
Acknowledgements We are grateful to Mr. Chengzhong Zhang for the help during the initial phase of this project and anonymous reviewers for the very valuable comments that helped improve this paper.
5